OpenStudy (anonymous):

I dont understand this Using the discriminant how many real number solutions does this equation have 3x^2-2=5x

OpenStudy (er.mohd.amir):

ax2+bx+c=0 discriminate=b*b-4ac find it

OpenStudy (er.mohd.amir):

what is value

OpenStudy (anonymous):

So I would set it up how

OpenStudy (er.mohd.amir):

a=3,b=-5,c=-2 so D=-5*-5-4*3*-2=25+24=49 taking sqrt give 7 and -7 so two real solution because 7 and -7

OpenStudy (er.mohd.amir):

3x^2-5x-2=0 compare with ax^2+bx+c=0 give value of a,b,c.

OpenStudy (anonymous):

So you solve it normally and take the square root

OpenStudy (anonymous):

3 -5 2

OpenStudy (er.mohd.amir):

since D*D=b*b-4*a*c if this value is -ve then U can not find sqrt of D then no real solution .

OpenStudy (er.mohd.amir):

it give no. of solution not solution of x.

OpenStudy (anonymous):

Oh ok that makes some sense thank you for helping me